# FuseMoE: Mixture-of-Experts Transformers for Fleximodal Fusion

### v2

@@ I.3 Over-specified Setting
In this appendix, we leverage proof techniques in Appendix K.1 to derive the density estimation rate under the over-specified setting in Theorem 3.1. Recall that under this setting, the MLE $\widehat{G}_{n}$ belongs to $\mathcal{G}_{k}(\Omega)$, i.e. the set of all mixing measures with at most $k>k_{*}$ components, where $k$ is unknown.
It is worth noting that if we can adapt the result of part (i) of Lemma K.1 to the over-specified settings, then other results presented in Appendix K.1 will also hold true. Therefore, our main goal is to establish following bound for the covering number of the metric space $(\mathcal{P}_{k}(\Omega),\|\cdot\|_{1})$:
$\log N(\eta,\mathcal{P}_{k}(\Omega),\|\cdot\|_{1})\lesssim\log(1/\eta),$
for any $20<\eta<1/$ 2, where $\mathcal{P}_{k}(\Omega):=\{\overline{p}_{G}(Y|X):G\in\mathcal{G}_{k}(\Omega)\}$.
To facilitate the presentation, we reuse the notations defined in Appendix K.1 with $\mathcal{G}_{k_{*}}(\Omega)$ being substituted by $\mathcal{G}_{k}(\Omega)$. Next, let us recall other essential notations for this proof.
Firstly, we define $\Theta=\{(a,b,\nu)\in\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}_{+}:(\beta,W,a,b,\nu)\in\Omega\}$, and $\overline{\Theta}_{\eta}$ is an $\eta$-cover of $\Theta$. Additionally, we also denote $\Delta:=\{(\beta,W)\in\mathbb{R}\times\mathbb{R}^{d}:(\beta,W,a,b,\nu)\in\Omega\}$, and $\overline{\Delta}_{\eta}$ be an $\eta$-cover of $\Delta$. Next, for each mixing measure $G=\sum_{i=1}^{k}\exp(\beta_{i})\delta_{(W_{i},a_{i},b_{i},\nu_{i})}\in\mathcal{G}_{k}(\Omega)$, we denote $G^{\prime}=\sum_{i=1}^{k}\exp(\beta_{i})\delta_{({W}_{i},\overline{a}_{i},\overline{b}_{i},\overline{\nu}_{i})}$ in which $(\overline{a}_{i},\overline{b}_{i},\overline{\nu}_{i})\in\overline{\Theta}_{\eta}$ is the closest point to $(a_{i},b_{i},\nu_{i})$ in this set for any $i\in[k]$. We also consider another mixing measure $\overline{G}:=\sum_{i=1}^{k}\exp(\overline{\beta}_{i})\delta_{(\overline{W}_{i},\overline{a}_{i},\overline{b}_{i},\overline{\nu}_{i})}\in\mathcal{G}_{k}(\Omega)$ where $(\overline{\beta}_{i},\overline{W}_{i})\in\overline{\Delta}_{\eta}$ is the closest point to $(\beta_{i},W_{i})$ in this set for any $i\in[k]$.
Subsequently, we define
:=$\mathcal{L}:=\{\overline{p}_{\overline{G}}\in\mathcal{P}_{k}(\Omega):(\overline{\beta}_{i},\overline{W}_{i})\in\overline{\Delta}_{\eta},\ (\overline{a}_{i},\overline{b}_{i},\overline{\nu}_{i})\in\overline{\Theta}_{\eta}\}.$
We demonstrate that $\mathcal{L}$ is an $\eta$-cover of the metric space $(\mathcal{P}_{k}(\Omega),\|\cdot\|_{1})$, that is, for any $\overline{p}_{G}\in\mathcal{P}_{k}(\Omega)$, there exists a density $\overline{p}_{\overline{G}}\in\mathcal{L}$ such that $\|\overline{p}_{G}-\overline{p}_{\overline{G}}\|_{1}\leq\eta$. By the triangle inequality, we have
$\|\overline{p}_{G}-\overline{p}_{\overline{G}}\|_{1}\leq\|\overline{p}_{G}-\overline{p}_{G^{\prime}}\|_{1}+\|\overline{p}_{G^{\prime}}-\overline{p}_{\overline{G}}\|_{1}.$ (23)
From the formulation of $G^{\prime}$, we get that
$\|\overline{p}_{G}-\overline{p}_{G^{\prime}}\|_{1}$ $\leq\sum_{i=1}^{k}\int_{\mathcal{X}\times\mathcal{Y}}\Big{|}f(Y|a_{i}^{\top}X+b_{i},\nu_{i})-f(Y|\overline{a}_{i}^{\top}X+\overline{b}_{i},\overline{\nu}_{i})\Big{|}\mathrm{d}(X,Y)$
$\lesssim\sum_{i=1}^{k}\int_{\mathcal{X}\times\mathcal{Y}}\Big{(}\|a_{i}-\overline{a}_{i}\|+|b_{i}-\overline{b}_{i}|+|\nu_{i}-\overline{\nu}_{i}|\Big{)}\mathrm{d}(X,Y)$
$\lesssim\eta$ (24)
Based on inequalities in equations equation 23 and equation 24, it is sufficient to show that $\|\overline{p}_{G^{\prime}}-\overline{p}_{\overline{G}}\|_{1}\lesssim\eta$. For any $\overline{\tau}\in[\overline{q}]$, let us define
$\overline{\mathcal{X}}_{\overline{\tau}}$ :=$:=\{x\in\mathcal{X}:-\|W_{i}-x\|\geq-\|W_{i^{\prime}}-x\|,\ \forall i\in\{\overline{\tau}_{1},\ldots,\overline{\tau}_{\overline{K}}\},i^{\prime}\in\{\overline{\tau}_{\overline{K}+1},\ldots,\overline{\tau}_{k}\}\},$
$\mathcal{X}^{\prime}_{\overline{\tau}}$ :=$:=\{x\in\mathcal{X}:-\|\overline{W}_{i}-x\|\geq-\|\overline{W}_{i^{\prime}}-x\|,\ \forall i\in\{\overline{\tau}_{1},\ldots,\overline{\tau}_{\overline{K}}\},i^{\prime}\in\{\overline{\tau}_{\overline{K}+1},\ldots,\overline{\tau}_{k}\}\}.$
Since the $\mathrm{softmax}$ function is differentiable, it is a Lipschitz function with some Lipschitz constant $L\geq 0$. Additionally, let us denote
$\pi_{\overline{\tau}}(X):=\Big{(}-\|W_{\overline{\tau}_{i}}-x\|+\beta^{\top}_{\overline{\tau}_{i}}\Big{)}_{i=1}^{\overline{K}};\qquad\overline{\pi}_{\overline{\tau}}(X):=\Big{(}-\|\overline{W}_{\overline{\tau}_{i}}-x\|+\overline{\beta}^{\top}_{\overline{\tau}_{i}}\Big{)}_{i=1}^{\overline{K}},$
for any $\overline{K}$-element subset $\{\overline{\tau}_{1},\ldots\overline{\tau}_{K}\}$ of $\{1,\ldots,k\}$. Then, we have
$\|\mathrm{softmax}(\pi_{\overline{\tau}}(X))-\mathrm{softmax}(\overline{\pi}_{\overline{\tau}}(X))\|$ $\leq L\cdot\|\pi_{\overline{\tau}}(X)-\overline{\pi}_{\overline{\tau}}(X)\|$
$\leq L\cdot\sum_{i=1}^{\overline{K}}\Big{(}\Big{|}\|W_{\overline{\tau}_{i}}-X\|-\|\overline{W}_{\overline{\tau}_{i}}-X\|\Big{|}+|\beta_{\overline{\tau}_{i}}-\overline{\beta}_{\overline{\tau}_{i}}|\Big{)}$
$\leq L\cdot\sum_{i=1}^{\overline{K}}\Big{(}\|W_{\overline{\tau}_{i}}-\overline{W}_{\overline{\tau}_{i}}\|+\beta_{\overline{\tau}_{i}}-\overline{\beta}_{\overline{\tau}_{i}}|\Big{)}$
$\leq L\cdot\sum_{i=1}^{\overline{K}}\Big{(}\eta+\eta\Big{)}$
$\lesssim\eta.$
By arguing similarly to the proof of Lemma L.2 in Appendix L, we receive that either $\overline{\mathcal{X}}_{\overline{\tau}}=\mathcal{X}^{\prime}_{\overline{\tau}}$ or $\overline{\mathcal{X}}_{\overline{\tau}}$ has measure zero for any $\overline{\tau}\in[\overline{q}]$. As a result, we deduce that
$\|\overline{p}_{G^{\prime}}-\overline{p}_{G_{*}}\|_{1}$
$\leq\sum_{\overline{\tau}=1}^{\overline{q}}\int_{\overline{\mathcal{X}}_{\overline{\tau}}\times\mathcal{Y}}|\overline{p}_{G^{\prime}}(Y|X)-\overline{p}_{\overline{G}}(Y|X)|\mathrm{d}(X,Y)$
$\leq\sum_{\overline{\tau}=1}^{\overline{q}}\int_{\overline{\mathcal{X}}_{\overline{\tau}}\times\mathcal{Y}}\sum_{i=1}^{\overline{K}}\Big{|}\mathrm{softmax}(\pi_{\overline{\tau}}(X)_{i})-\mathrm{softmax}(\overline{\pi}_{\overline{\tau}}(X)_{i})\Big{|}\cdot\Big{|}f(Y|\overline{a}_{\overline{\tau}_{i}}^{\top}X+\overline{b}_{\overline{\tau}_{i}},\overline{\nu}_{\overline{\tau}_{i}})\Big{|}\mathrm{d}(X,Y)$
$\lesssim\eta.$
Thus, $\mathcal{L}$ is an $\eta$-cover of the metric space $(\mathcal{P}_{k}(\Omega),\|\cdot\|_{1})$, which implies that
$N(\eta,\mathcal{P}_{k}(\Omega),\|\cdot\|_{1})\lesssim|\overline{\Delta}_{\eta}|\times|\overline{\Theta}_{\eta}|\leq\mathcal{O}(\eta^{-(d+1)k})\times\mathcal{O}(\eta^{-(d+2)k})=\mathcal{O}(\eta^{-(2d+3)k}).$ (25)
Hence, we reach the conclusion that $\log N(\eta,\mathcal{P}_{k}(\Omega),\|\cdot\|_{1})\lesssim\log(1/\eta)$.
K.4 Proof of Theorem 3.2
@@
$\mathbb{E}_{X}[V(\overline{p}_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]\gtrsim\mathcal{D}_{2}(G,G_{*}),$
@@
  • Inequality A. $\inf_{G\in\mathcal{G}_{k}(\Omega):\mathcal{D}_{2}(G,G_{*})\leq\varepsilon^{\prime}}\dfrac{\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]}{\mathcal{D}_{2}(G,G_{*})}>0$;
  • Inequality B. $\inf_{G\in\mathcal{G}_{k}(\Omega):\mathcal{D}_{2}(G,G_{*})>\varepsilon^{\prime}}\dfrac{\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]}{\mathcal{D}_{2}(G,G_{*})}>0$,
@@
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{G}_{k}(\Theta):\mathcal{D}_{2}(G,G_{*})\leq\varepsilon}\frac{\mathbb{E}_{X}[V(\overline{p}_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]}{\mathcal{D}_{2}(G,G_{*})}>0.$ (26)
@@
Assume that the above claim does not hold true, then there exists a sequence of mixing measures $G_{n}:=\sum_{i=1}^{k_{n}}\exp(\beta_{i}^{n})\delta_{(W_{i}^{n},a_{i}^{n},b_{i}^{n},\nu_{i}^{n})}\in\mathcal{G}_{k}(\Omega)$ such that both $\mathcal{D}_{2}(G_{n},G_{*})$ and $\mathbb{E}_{X}[V(\overline{p}_{G_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]/\mathcal{D}_{2}(G_{n},G_{*})$ go to zero as $n\to\infty$. Since $\mathcal{D}_{2}(G_{n},G_{*})\to 0$, we deduce that $\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i}^{n})\to\exp(\beta_{j}^{*})$ and $(W_{i}^{n},a_{i}^{n},b_{i}^{n},\nu_{i}^{n})\to(W_{j}^{*},a_{j}^{*},b_{j}^{*},\nu_{j}^{*})$ for all $i\in\mathcal{A}_{j}$ and $j\in[k_{*}]$. WLOG, we may assume that
$\mathcal{D}_{2}(G_{n},G_{*})=\sum_{\begin{subarray}{c}j\in[K],\\ |\mathcal{A}_{j}|>1\end{subarray}}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i}^{n})\Big{[}\|\Delta W_{ij}^{n}\|^{2}+\|\Delta a_{ij}^{n}\|^{2}+|\Delta b_{ij}^{n}|^{\bar{r}_{j}}+|\Delta\nu_{ij}^{n}|^{\frac{\bar{r}_{j}}{2}}\Big{]}$
$+\sum_{\begin{subarray}{c}j\in[K],\\ |\mathcal{A}_{j}|=1\end{subarray}}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i}^{n})\Big{[}\|\Delta W_{ij}^{n}\|+\|\Delta a_{ij}^{n}\|+|\Delta b_{ij}^{n}|+|\Delta\nu_{ij}^{n}|\Big{]}+\sum_{j=1}^{K}\Big{|}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i}^{n})-\exp(\beta_{j}^{*})\Big{|}.$ (27)
Regarding the top-$K$ selection in the conditional density $p_{G_{*}}$, we partition the covariate space $\mathcal{X}$ in a similar fashion to Appendix K.2. More specifically, we consider $q=\binom{k_{*}}{K}$ subsets $\{\tau_{1},\ldots,\tau_{K}\}$ of $\{1,\ldots,k_{*}\}$ for any $\tau\in[q]$, and denote $\{\tau_{K+1},\ldots,\tau_{k_{*}}\}:=[k_{*}]\setminus\{\tau_{1},\ldots,\tau_{K}\}$. Then, we define
$\mathcal{X}^{*}_{\tau}:=\Big{\{}x\in\mathcal{X}:-\|W_{j}^{*}-x\|\geq-\|W^{*}_{j^{\prime}}-x\|,\forall j\in\{\tau_{1},\ldots,\tau_{K}\},j^{\prime}\in\{\tau_{K+1},\ldots,\tau_{k_{*}}\}\Big{\}},$
for any $\tau\in[q]$. On the other hand, we need to introduce a new partition method of the covariate space for the weight selection in the conditional density $p_{G_{n}}$. In particular, let $\overline{K}\in\mathbb{N}$ such that $\max_{\{\tau_{j}\}_{j=1}^{K}\subset[k_{*}]}\sum_{j=1}^{K}|\mathcal{A}_{\tau_{j}}|\leq\overline{K}\leq k$ and $\overline{q}:=\binom{k}{\overline{K}}$. Then, for any $\overline{\tau}\in[\overline{q}]$, we denote $(\overline{\tau}_{1},\ldots,\overline{\tau}_{k})$ as a subset of $[k]$ and $\{\overline{\tau}_{\overline{K}+1},\ldots,\overline{\tau}_{k}\}:=[k]\setminus\{\overline{\tau}_{1},\ldots,\overline{\tau}_{\overline{K}}\}$. Additionally, we define
$\mathcal{X}^{n}_{\overline{\tau}}:=\Big{\{}x\in\mathcal{X}:-\|W_{i}^{n}-x\|\geq-\|W^{n}_{i^{\prime}}-x\|,\forall i\in\{\overline{\tau}_{1},\ldots,\overline{\tau}_{\overline{K}}\},i^{\prime}\in\{\overline{\tau}_{\overline{K}+1},\ldots,\overline{\tau}_{k}\}\Big{\}}.$
Let $X\in\mathcal{X}^{*}_{\tau}$ for some $\tau\in[q]$ such that $\{\tau_{1},\ldots,\tau_{K}\}=\{1,\ldots,K\}$. If $\{\overline{\tau}_{1},\ldots\overline{\tau}_{\overline{K}}\}\neq\mathcal{A}_{1}\cup\ldots\cup\mathcal{A}_{K}$ for any $\overline{\tau}\in[\overline{q}]$, then $\mathbb{E}_{X}[V(\overline{p}_{G_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]/\mathcal{D}_{2}(G_{n},G_{*})\not\to 0$ as $n$ tends to infinity. This contradicts the fact that this term must approach zero. Therefore, we only need to consider the scenario when there exists $\overline{\tau}\in[\overline{q}]$ such that $\{\overline{\tau}_{1},\ldots\overline{\tau}_{\overline{K}}\}=\mathcal{A}_{1}\cup\ldots\cup\mathcal{A}_{K}$. Recall that we have $(W_{i}^{n},\beta_{i}^{n})\to(W_{j}^{*},\beta_{j}^{*})$ as $n\to\infty$ for any $j\in[k_{*}]$ and $i\in\mathcal{A}_{j}$. Thus, for any arbitrarily small $\eta_{j}>0$, we have that $\|W_{i}^{n}-W_{j}^{*}\|\leq\eta_{j}$ and $|\beta_{i}^{n}-\beta_{j}^{*}|\leq\eta_{j}$ for sufficiently large $n$. Then, it follows from Lemma L.2 that $\mathcal{X}^{*}_{\tau}=\mathcal{X}^{n}_{\overline{\tau}}$ for sufficiently large $n$. This result indicates that $X\in\mathcal{X}^{n}_{\overline{\tau}}$.
Then, we can represent the conditional densities $p_{G_{*}}(Y|X)$ and $p_{G_{n}}(Y|X)$ for any sufficiently large $n$ as follows:
$p_{G_{*}}(Y|X)$ $=\sum_{j=1}^{K}\frac{\exp(-\|W_{j}^{*}-X\|+\beta_{j}^{*})}{\sum_{j^{\prime}=1}^{K}\exp(-\|W^{*}_{j^{\prime}}-x\|+\beta^{*}_{0j^{\prime}})}\cdot f(Y|(a_{j}^{*})^{\top}X+b_{j}^{*},\nu_{j}^{*}),$
$\overline{p}_{G_{n}}(Y|X)$ $=\sum_{j=1}^{K}\sum_{i\in\mathcal{A}_{j}}\frac{\exp(-\|W_{i}^{n}-x\|+\beta_{i}^{n})}{\sum_{j^{\prime}=1}^{K}\sum_{i^{\prime}\in\mathcal{A}_{j^{\prime}}}\exp(-\|W^{n}_{i^{\prime}}-x\|+\beta^{n}_{0i^{\prime}})}\cdot f(Y|(a_{i}^{n})^{\top}X+b_{i}^{n},\nu_{i}^{n}).$
@@
$Q_{n}$ $=\sum_{j=1}^{K}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i}^{n})\Big{[}F(Y|X;W_{i}^{n},a_{i}^{n},b_{i}^{n},\nu_{i}^{n})-F(Y|X;W_{j}^{*},a_{j}^{*},b_{j}^{*},\nu_{j}^{*})\Big{]}$
$-\sum_{j=1}^{K}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i}^{n})\Big{[}H(Y|X;W_{i}^{n})-H(Y|X;W_{j}^{*})\Big{]}$
$+\sum_{j=1}^{K}\Big{[}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i}^{n})-\exp(\beta_{j}^{*})\Big{]}\Big{[}F(Y|X;W_{j}^{*},a_{j}^{*},b_{j}^{*},\nu_{j}^{*})-H(Y|X,W_{j}^{*})\Big{]}$
@@
$Q_{n}$ $=\sum_{j=1}^{K}\sum_{|\alpha_{1}|=0}^{\bar{r}_{j}}\sum_{|\alpha_{2}|=0}^{\bar{r}_{j}-|\alpha_{1}|}\sum_{\eta=0}^{2(\bar{r}_{j}-|\alpha_{1}|-|\alpha_{2}|)}S^{n}_{j,\alpha_{1},\alpha_{2},\eta}\cdot X^{\alpha_{2}}\cdot\dfrac{\partial^{|\alpha_{1}|}g}{\partial W^{\alpha_{1}}}(X;W_{j}^{*})\cdot\frac{\partial^{|\alpha_{2}|+\eta}f}{\partial h_{1}^{|\alpha_{2}|+\eta}}(Y|(a_{j}^{*})^{\top}X+b_{j}^{*},\nu_{j}^{*})$
$+\sum_{j=1}^{K}\sum_{|\gamma|=0}^{1+\mathbf{1}_{\{|\mathcal{A}_{j}|>1\}}}T^{n}_{j,\gamma}\cdot\dfrac{\partial^{|\gamma|}g}{\partial W^{\gamma}}(X;W_{j}^{*})p_{G_{n}}(Y|X)+R_{3}(X,Y)+R_{4}(X,Y)+R_{5}(X,Y)+R_{6}(X,Y).$ (28)
@@
$\frac{1}{\mathcal{D}_{2}(G_{n},G_{*})}\cdot\Big{[}\sum_{j=1}^{K}$ $\Big{|}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i}^{n})-\exp(\beta_{j}^{*})\Big{|}$
@@
This result implies that we can find some index $j^{\prime}\in[K]:|\mathcal{A}_{j^{\prime}}|>1$ that satisfies
@@
Recall that $\mathbb{E}_{X}[V(\overline{p}_{G_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]/\mathcal{D}_{2}(G_{n},G_{*})\to 0$ as $n\to\infty$. Then, by applying the Fatou’s lemma, we get
$0=\lim_{n\to\infty}\dfrac{\mathbb{E}_{X}[V(\overline{p}_{G_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]}{\mathcal{D}_{2}(G_{n},G_{*})}\geq\frac{1}{2}\cdot\int\liminf_{n\to\infty}\frac{|\overline{p}_{G_{n}}(Y|X)-p_{G_{*}}(Y|X)|}{\mathcal{D}_{2}(G_{n},G_{*})}\mathrm{d}(X,Y),$
which implies that $|\overline{p}_{G_{n}}(Y|X)-p_{G_{*}}(Y|X)|/\mathcal{D}_{2}(G_{n},G_{*})\to 0$ as $n\to\infty$ for almost surely $(X,Y)$.
@@
$\sum_{j=1}^{K}\sum_{|\alpha_{1}|=0}^{\bar{r}_{j}}\sum_{|\alpha_{2}|=0}^{\bar{r}_{j}-|\alpha_{1}|}\sum_{\eta=0}^{2(\bar{r}_{j}-|\alpha_{1}|-|\alpha_{2}|)}\xi_{j,\alpha_{1},\alpha_{2},\eta}\cdot X^{\alpha_{2}}\cdot\dfrac{\partial^{|\alpha_{1}|}g}{\partial W^{\alpha_{1}}}(X;W_{j}^{*})\cdot\frac{\partial^{|\alpha_{2}|+\eta}f}{\partial h_{1}^{|\alpha_{2}|+\eta}}(Y|(a_{j}^{*})^{\top}X+b_{j}^{*},\nu_{j}^{*})$

### v4

@@ I.3 Over-specified Setting
$\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]\gtrsim\mathcal{D}_{2}(G,G_{*}),$
@@
  • Inequality A. $\inf_{G\in\mathcal{G}_{k}(\Theta):\mathcal{D}_{2}(G,G_{*})\leq\varepsilon^{\prime}}\dfrac{\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]}{\mathcal{D}_{2}(G,G_{*})}>0$;
  • Inequality B. $\inf_{G\in\mathcal{G}_{k}(\Theta):\mathcal{D}_{2}(G,G_{*})>\varepsilon^{\prime}}\dfrac{\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]}{\mathcal{D}_{2}(G,G_{*})}>0$,
@@
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{G}_{k}(\Theta):\mathcal{D}_{2}(G,G_{*})\leq\varepsilon}\frac{\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]}{\mathcal{D}_{2}(G,G_{*})}>0.$ (22)
@@
Assume that the above claim does not hold true, then there exists a sequence of mixing measures $G_{n}:=\sum_{i=1}^{k_{n}}\exp(\beta_{i}^{n})\delta_{(W_{i}^{n},a_{i}^{n},b_{i}^{n},\nu_{i}^{n})}\in\mathcal{G}_{k}(\Theta)$ such that both the terms $\mathcal{D}_{2}(G_{n},G_{*})$ and $\mathbb{E}_{X}[V(p_{G_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]/\mathcal{D}_{2}(G_{n},G_{*})$ go to zero as $n\to\infty$. Let us recall the formulation of the loss $\mathcal{D}_{2}(G_{n},G_{*})$:
$\mathcal{D}_{2}(G_{n},G_{*})=\sum_{\begin{subarray}{c}j\in[k_{*}],\\ |\mathcal{A}_{j}|>1\end{subarray}}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i}^{n})\Big{[}\|\Delta W_{ij}^{n}\|^{2}+\|\Delta a_{ij}^{n}\|^{2}+|\Delta b_{ij}^{n}|^{\bar{r}_{j}}+|\Delta\nu_{ij}^{n}|^{\frac{\bar{r}_{j}}{2}}\Big{]}$
$+\sum_{\begin{subarray}{c}j\in[k_{*}],\\ |\mathcal{A}_{j}|=1\end{subarray}}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i}^{n})\Big{[}\|\Delta W_{ij}^{n}\|+\|\Delta a_{ij}^{n}\|+|\Delta b_{ij}^{n}|+|\Delta\nu_{ij}^{n}|\Big{]}+\sum_{j=1}^{k_{*}}\Big{|}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i}^{n})-\exp(\beta_{j}^{*})\Big{|}.$ (23)
Since $\mathcal{D}_{2}(G_{n},G_{*})\to 0$, we deduce that $\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i}^{n})\to\exp(\beta_{j}^{*})$ and $(W_{i}^{n},a_{i}^{n},b_{i}^{n},\nu_{i}^{n})\to(W_{j}^{*},a_{j}^{*},b_{j}^{*},\nu_{j}^{*})$ for all $i\in\mathcal{A}_{j}$ and $j\in[k_{*}]$.
@@
$Q_{n}$ $=\sum_{j=1}^{k_{*}}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i}^{n})\Big{[}F(Y|X;W_{i}^{n},a_{i}^{n},b_{i}^{n},\nu_{i}^{n})-F(Y|X;W_{j}^{*},a_{j}^{*},b_{j}^{*},\nu_{j}^{*})\Big{]}$
$-\sum_{j=1}^{k_{*}}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i}^{n})\Big{[}H(Y|X;W_{i}^{n})-H(Y|X;W_{j}^{*})\Big{]}$
$+\sum_{j=1}^{k_{*}}\Big{[}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i}^{n})-\exp(\beta_{j}^{*})\Big{]}\Big{[}F(Y|X;W_{j}^{*},a_{j}^{*},b_{j}^{*},\nu_{j}^{*})-H(Y|X,W_{j}^{*})\Big{]}$
@@
$Q_{n}$ $=\sum_{j=1}^{k_{*}}\sum_{|\alpha_{1}|=0}^{\bar{r}_{j}}\sum_{|\alpha_{2}|=0}^{\bar{r}_{j}-|\alpha_{1}|}\sum_{\eta=0}^{2(\bar{r}_{j}-|\alpha_{1}|-|\alpha_{2}|)}S^{n}_{j,\alpha_{1},\alpha_{2},\eta}\cdot X^{\alpha_{2}}\cdot\dfrac{\partial^{|\alpha_{1}|}g}{\partial W^{\alpha_{1}}}(X;W_{j}^{*})\cdot\frac{\partial^{|\alpha_{2}|+\eta}f}{\partial h_{1}^{|\alpha_{2}|+\eta}}(Y|(a_{j}^{*})^{\top}X+b_{j}^{*},\nu_{j}^{*})$
$+\sum_{j=1}^{k_{*}}\sum_{|\gamma|=0}^{1+\mathbf{1}_{\{|\mathcal{A}_{j}|>1\}}}T^{n}_{j,\gamma}\cdot\dfrac{\partial^{|\gamma|}g}{\partial W^{\gamma}}(X;W_{j}^{*})p_{G_{n}}(Y|X)+R_{3}(X,Y)+R_{4}(X,Y)+R_{5}(X,Y)+R_{6}(X,Y).$ (24)
@@
$\frac{1}{\mathcal{D}_{2}(G_{n},G_{*})}\cdot\Big{[}\sum_{j=1}^{k_{*}}$ $\Big{|}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i}^{n})-\exp(\beta_{j}^{*})\Big{|}$
@@
This result implies that we can find some index $j^{\prime}\in[k_{*}]:|\mathcal{A}_{j^{\prime}}|>1$ that satisfies
@@
Recall that $\mathbb{E}_{X}[V(p_{G_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]/\mathcal{D}_{2}(G_{n},G_{*})\to 0$ as $n\to\infty$. Then, by applying the Fatou’s lemma, we get
$0=\lim_{n\to\infty}\dfrac{\mathbb{E}_{X}[V(p_{G_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]}{\mathcal{D}_{2}(G_{n},G_{*})}\geq\frac{1}{2}\cdot\int\liminf_{n\to\infty}\frac{|p_{G_{n}}(Y|X)-p_{G_{*}}(Y|X)|}{\mathcal{D}_{2}(G_{n},G_{*})}\mathrm{d}(X,Y),$
which implies that $|p_{G_{n}}(Y|X)-p_{G_{*}}(Y|X)|/\mathcal{D}_{2}(G_{n},G_{*})\to 0$ as $n\to\infty$ for almost surely $(X,Y)$.
@@
$\sum_{j=1}^{k_{*}}\sum_{|\alpha_{1}|=0}^{\bar{r}_{j}}\sum_{|\alpha_{2}|=0}^{\bar{r}_{j}-|\alpha_{1}|}\sum_{\eta=0}^{2(\bar{r}_{j}-|\alpha_{1}|-|\alpha_{2}|)}\xi_{j,\alpha_{1},\alpha_{2},\eta}\cdot X^{\alpha_{2}}\cdot\dfrac{\partial^{|\alpha_{1}|}g}{\partial W^{\alpha_{1}}}(X;W_{j}^{*})\cdot\frac{\partial^{|\alpha_{2}|+\eta}f}{\partial h_{1}^{|\alpha_{2}|+\eta}}(Y|(a_{j}^{*})^{\top}X+b_{j}^{*},\nu_{j}^{*})$